There is a 6M stock solution of “Solution 2” provided. Develop a procedure and calculate the necessary volumes to dilute the stock solution to make 125 mL of a 0.48M solution.

Volume of solution 2:

Volume of Water:
Answer #1

stock solution                           dilute solution

M1 = 6M                                  M2   = 0.48M

V1 =                                          V2 = 125ml

           M1V1    = M2V2

             V1    = M2V2/M1

                     = 0.48*125/6  

                      = 10ml

volume of solution 2 : 10ml

volume of water added = V2-V1

                                       = 125-10

                                         = 115ml >>>>>answer